Skin color in frogs is determined by two alleles of gene B (BB and Bb=brown, bb=green). After a drought, 20 frogs remain in the pond. Their genotypes are: 9 BB, 10 Bb, 1 bb. If Hardy-Weinberg equilibrium is established in this new population, what would be the frequency of green frogs in the next generation? A. 0.04 B. 0.09 C. 0.3
Answer #1

Answer:B. 0.09

Explanation:

9 BB, 10 Bb, 1 bb

p^2=9/20=0.45

q^2=1/20 = 0.05

2pq=10 = 10/20 = 0.5

p=0.45+0.25 =0.7

q= 0.05 + 0.25 = 0.3

the frequency of green frogs in the next generation

q^2= 0.3 x 0.3 = 0.09